# Robust Offline Imitation Learning from Diverse Auxiliary Data

### v1

@@ 4.3 Reward-weighted behavioral cloning
$\min_{\pi}\underset{(s,a)\sim{\mathcal{D}}_{E}}{\mathbb{E}}\left[-\log\pi(a|s)\right]+\\ \alpha\underset{(s,a)\sim{\mathcal{D}}_{O}}{\mathbb{E}}\left[-\log\pi(a|s)\cdot\tilde{r}(s,a)\cdot\mathbbm{1}[\tilde{r}>\tau]\right]\ .$ (3)
@@ 4.4 TD learning using learned rewards
Despite not meeting the reward threshold, sub-optimal state-action tuples possess the potential to enhance the learned policy, augmenting its robustness against distribution shifts during deployment. This is due to the broader coverage of the state space within the auxiliary dataset, surpassing the limited span of the small expert dataset. In the absence of access to expert behavior in these states for direct learning, we employ a shortest path strategy to leverage these samples. More precisely, our objective on these states is to steer the policy toward states observed by the expert and subsequently imitate the expert’s behavior accordingly. To accomplish this, we use a TD3-style [9] algorithm to learn a Q-function using the approximated rewards and then direct the policy to maximize the long-term return on these expert-unobserved states. The Q-function is learned as follows,
@@
$\min_{\pi}\underset{(s,a)\sim{\mathcal{D}}_{E}}{\mathbb{E}}\left[-\log\pi(a|s)\right]+\\ \alpha\underset{(s,a)\sim{\mathcal{D}}_{O}}{\mathbb{E}}\left[-\log\pi(a|s)\cdot\tilde{r}(s,a)\cdot\mathbbm{1}[\tilde{r}>\tau]\right]+\\ \beta\underset{s\sim\mathcal{D}_{E}\cup\mathcal{D}_{O}}{\mathbb{E}}\left[-Q(s,\pi(s))\right]\ .$ (6)
Here, $\beta$ is a hyperparameter which controls the contribution of the Q-loss. By maximizing this Q-function alongside the weighted BC objective, we incentivize the policy to: 1) act optimally in states where we have expert actions, and 2) guide the agent efficiently from expert-unobserved states to expert-observed states and act optimally subsequently. The pseudo-code of our algorithm is presented in Appendix A.
@@ 5 Experiments
$\bullet$ How robust is ROIDA when the quality of the auxiliary data varies? (Sec. 5.3.1)
$\bullet$ How does ROIDA compare to other methods as the size of the expert dataset changes? (Sec. 5.3.2)
@@ 5.1 Experimental setup
We conduct experiments on locomotion and manipulation tasks from the D4RL benchmark [7].
Locomotion. We use 4 MuJoCo environments for the locomotion tasks: hopper, halfcheetah, walker2d, and ant. For expert demonstrations in each environment, we use the corresponding dataset: hopper-expert-v2, halfcheetah-expert-v2, walker2d-expert-v2, and ant-expert-v2. For the sub-optimal demonstrations, we source trajectories from the respective random-v2 datasets. To create $\mathcal{D}_{E}$, we randomly sample 3, 5, or 7 trajectories per environment depending on the chosen setting. For $\mathcal{D}_{O}$, we create 3 settings per environment: 1000 randomly sampled sub-optimal trajectories plus another 0, 3, or 5 expert trajectories. This allows us to test our method’s ability to identify and leverage expert demonstrations within different mixes of sub-optimal and expert data.
@@
Manipulation. We evaluate our method on 4 ADROIT manipulation tasks from using a simulated 24 DoF hand: pen twirling, hammering a nail, opening a door, and relocating a ball. For expert demonstrations, we sample 50 trajectories from pen-expert-v1, hammer-expert-v1, door-expert-v1, and relocate-expert-v1 respectively. For the sub-optimal demonstrations, we use 1000 trajectories from the datasets pen-cloned-v1, hammer-cloned-v1, door-cloned-v1, and relocate-cloned-v1, plus 0, 30, or 50 expert trajectories from the corresponding expert-v1 datasets.
Note that for all the evaluations, we report the average of the mean normalized score for the last 10 evaluations of training over 5 random seeds. Additional implementation details for our method can be found in Appendix D. We also present evaluations using the rliable [2] framework in Appendix F.
@@
5.2 Baselines
We compare ROIDA against the following algorithms:
$\bullet$ BC-exp: BC-exp denotes behavioral cloning solely on the expert dataset $\mathcal{D}_{E}$. Since $\mathcal{D}_{E}$ contains only a small number of expert demonstrations, training only on this data can lead to degraded performance at test time due to compounding errors caused by distribution shift.
$\bullet$ BC-all: In this case, both $\mathcal{D}_{E}$ and $\mathcal{D}_{O}$ are used to learn the policy. Despite the large number of demonstrations, a significant portion of them are random or low quality. As a result, the learned policy tends to be sub-optimal due to the inclusion of poor demonstrations.
$\bullet$ DWBC: DWBC [37] treats the auxiliary data as a mixture of expert and sub-optimal data, and utilizes PU learning to train a discriminator for weighted behavioral cloning. They perform a dual learning strategy where they alternately train the discriminator and the policy by taking the output of each model as an input to the other. Due to relying solely on BC, DWBC performs poorly when the number of expert transitions is low.
$\bullet$ DemoDICE: DemoDICE [13] conducts state-action distribution matching over $\mathcal{D}_{E}$ and introduces a regularization constraint to ensure the learned policy remains close to the behavior policy of $\mathcal{D}_{O}$. It shares the same drawbacks as DWBC due to the second term, resulting in a suboptimal policy especially when $\mathcal{D}_{O}$ contains a large proportion of noisy data.
$\bullet$ ORIL: ORIL [41] first learns a reward function and then performs Critic-Regularized Regression [35] to learn the policy by enriching the data using different augmentation strategies.
5.3 Results
5.3.1 Varying the quality of auxiliary data
Although DWBC achieves the second highest average performance, it degrades substantially compared to ROIDA when auxiliary data quality is poor (5/0 and 5/3). This stems from DWBC’s inability to extract useful information from poor quality unlabeled demos. In contrast, ROIDA can utilize the auxiliary data through temporal difference learning, despite poor quality. Both DemoDICE and ORIL perform worse in all settings.
Table 3: Ablation study on each module’s contribution to final policy performance. For each locomotion environment the performance is averaged across scenarios containing different number of expert trajectories in $\mathcal{D}_{O}$ ($5/0$, $5/3$, $/5$ 5).
Method Environment
Hopper HalfCheetah Walker2D Ant
ROIDA 86.58 p m 16.42 19.78 p m 6.58 108.68 p m 0.21 73.31 p m 15.79
- w/o reward-weighted BC $67.08$ p m 26.34 $14.19$ p m 5.70 $101.27$ p m 16.70 $60.25$ p m 17.14
- w/o modified reward $84.96$ p m 9.97 $17.61$ p m 6.55 $104.99$ p m 6.60 $62.81$ p m 13.95
5.3.2 Varying the size of the expert dataset
In Table 2, we study the effect of varying the size of the expert dataset on policy performance. We hold the auxiliary dataset fixed, with the number of expert trajectories in $\mathcal{D}_{O}$ set to 5, and vary the number of expert trajectories in $\mathcal{D}_{E}$ across 3, 5, and 7. ROIDA achieves comparable performance to DWBC, slightly outperforming in all 4 environments. As the amount of expert data in $\mathcal{D}_{E}$ increases, all the methods improve, as expected. In summary, ROIDA maintains competitive performance across varying expert dataset sizes. Due to space constraints, we present the results on the manipulation tasks in Appendix C.
5.4 Ablation study
In this section, we conduct ablation studies to validate the importance of each component in ROIDA, and also benchmark the performance as the scale of the auxiliary dataset is varied.
First, we systematically remove each part of ROIDA and evaluate the resulting performance impact. The ablation results in Table 3 present averaged results across three auxiliary dataset settings: 5/0, 5/3, 5/5 (denoting different auxiliary data qualities). The full ablation results per setting are presented in Appendix B.1.
We study the impact of ablating the reward weighted BC component by setting $\alpha=0$. Performance drops indicate the importance of filtering the data using a learned reward model. The precise reward assignment enables ROIDA to discern the varying quality of transitions in the data and utilize them accordingly.
We also study the impact of the reward function formulation by directly using the discriminator output as the reward. The results demonstrate that our reward formulation based on the DICE methods improves policy performance compared to directly using the discriminator rewards.
Second, we change the scale of the auxiliary dataset and analyze the impact on the results. We achieve this by including 25%, 50% or 100% of the entire random set from D4RL (instead of the fixed 1000 trajectories in the previous experiments). As shown in Fig. 3, even with an increased number of suboptimal demonstrations, performance improves as more expert trajectories are added. This clearly highlights ROIDA’s ability to distinguish expert data even when the ratio of expert to noisy data is highly skewed. We also observe that adding more auxiliary data while keeping the number of expert trajectories fixed slightly improves performance, indicating ROIDA’s ability to extract information from suboptimal data. We further analyze the impact of the reward threshold $\tau$ in Appendix B.2.
6 Conclusion
We propose ROIDA, a simple yet effective framework for offline imitation that can maximize utilization of an unlabeled auxiliary dataset of unknown quality alongside a small set of expert demonstrations. Unlike previous methods that make assumptions about auxiliary dataset quality, ROIDA can seamlessly leverage uncurated, unlabeled offline datasets without relying on any quality assumptions. We demonstrate ROIDA’s efficacy on multiple manipulation and locomotion tasks, encompassing a wide variety of auxiliary dataset quality settings. The consistent performance gains over baselines validate ROIDA’s ability to unlock the full potential of heterogeneous offline datasets without relying on quality assumptions. A valuable direction for future research is to adapt ROIDA to learn policies directly from multi-modal inputs on physical robots. We discuss the limitations and broader impacts of our method in Appendix G and Appendix H respectively.
Acknowledgement This work was partially supported by The National Science Foundation Award No. 2326309, National Institute for Food and Agriculture Award No. 2021-67022-33453 and the UC Multi-campus Research Programs Initiative.
Appendix A Pseudo-code
In this section, we present the pseudo-code of ROIDA in Algorithm 1.
2:  Train discriminator $d$ with non-negative PU learning following Equation 1
3:  for t=1 to T do
6:     Compute Q-function output values $Q(s,a)$ using sampled $(s,a)$ and $\mathcal{B}^{\pi}Q(s,a)$ using Equation 5
7:     Update $Q$ by minimizing the learning objective given in Equation 4
8:     if t mod $t_{freq}$ then
9:        Update $\pi$ by minimizing the learning objective in Equation 6
10:     end if
11:  end for
Appendix B Ablation study
B.1 Expanded ablation results
Table 4 shows the full ablation results for locomotion tasks across three auxiliary data settings: 5/0, 5/3, and 5/5. Removing any component from ROIDA decreases performance, with the most significant drop when ablating reward-weighted BC. This highlights the importance of filtering auxiliary data to leverage high-quality transitions. In summary, the ablation study validates the contribution of each proposed technique in ROIDA to effectively leverage diverse quality offline data.
Table 4: Ablation study on locomotion tasks as the number of expert demonstrations in $\mathcal{D}_{O}$ is increased. The best performing method on each task is highlighted in red and the second best in blue.
Environment Setting Method
ROIDA w/o reward-weighted BC w/o modified reward
Hopper 5 /  0 $84.63$ p m 16.01 $41.68$ p m 39.67 $78.07$ p m 11.00
5 /  3 $86.66$ p m 21.94 $78.06$ p m 16.17 $92.06$ p m 7.10
5 /  5 $88.45$ p m 8.46 $81.51$ p m 15.70 $84.73$ p m 11.26
Avg. 86.58 p m 16.42 67.08 p m 26.34 84.96 p m 9.97
HalfCheetah 5 /  0 $15.89$ p m 9.60 $9.42$ p m 1.48 $13.93$ p m 3.31
5 /  3 $18.73$ p m 3.67 $14.93$ p m 8.61 $16.35$ p m 4.44
5 /  5 $24.70$ p m 4.95 $18.22$ p m 4.61 $22.53$ p m 9.91
Avg. 19.78 p m 6.58 14.19 p m 5.70 17.61 p m 6.55
Walker2D 5 /  0 $108.73$ p m 0.28 $87.72$ p m 28.91 $104.65$ p m 7.79
5 /  3 $108.52$ p m 0.21 $108.11$ p m 0.66 $104.77$ p m 6.83
5 /  5 $108.794$ p m 0.09 $107.97$ p m 0.55 $105.56$ p m 4.81
Avg. 108.68 p m 0.21 101.27 p m 16.70 104.99 p m 6.60
Ant 5 /  0 $65.73$ p m 25.19 $57.42$ p m 20.89 $49.37$ p m 11.64
5 /  3 $77.52$ p m 4.98 $64.40$ p m 11.52 $62.94$ p m 10.27
5 /  5 $76.68$ p m 9.44 $58.95$ p m 17.68 $76.14$ p m 18.52
Avg. 73.31 p m 15.79 60.25 p m 17.14 62.81 p m 13.95
B.2 Impact of threshold
Our specific reward formulation leads to rewards in the range $[-2.19,2.19]$. Based on this, we decide to set the threshold $\tau=1$ to strike a balance between utilizing the “good” transitions in the auxiliary data and neglecting the poor quality transitions. To evaluate the impact of the threshold on the results, we vary on the Hopper environment as shown in Table 5.
Setting Reward threshold
5 /  0 $83.20$ p m 24.58 $84.63$ p m 16.01 $70.24$ p m 27.75
5 /  3 $86.82$ p m 6.85 $86.66$ p m 21.94 $81.91$ p m 18.33
5 /  5 $88.23$ p m 11.42 $88.45$ p m 8.46 $90.70$ p m 6.88
Appendix C Varying experts on manipulation tasks
Similar to locomotion experiment for increasing number of experts in $\mathcal{D}_{E}$, we vary the expert set for Adroit tasks in Table 6. We show the comparative performance of our model with 4 other baselines by changing the expert set from 50 to 70 while keeping the auxiliary data constant. Here, too, we observe improved average performance in our method compared to the other baselines.
Table 6: Imitation learning performance on Adroit tasks as the number of expert demonstrations in $\mathcal{D}_{E}$ is increased. The best performing method on each task is highlighted in red and the second best in blue.
Environment Setting Method
BC-exp BC-all DemoDICE ORIL DWBC ROIDA
@@ Appendix C Varying experts on manipulation tasks
Figure 4: rliable evaluation on the Locomotion benchmark.
Figure 5: rliable evaluation on the Adroit benchmark.
Appendix D Experimental details
D.1 Implementation

### v3

@@ 4.3 Reward-weighted behavioral cloning
$\min_{\pi}$ $\underset{(s,a)\sim{\mathcal{D}}_{E}}{\mathbb{E}}\left[-\log\pi(a|s)\right]+\alpha\underset{(s,a)\sim{\mathcal{D}}_{O}}{\mathbb{E}}\left[-\log\pi(a|s)\cdot\tilde{r}(s,a)\cdot\mathbbm{1}[\tilde{r}>\tau]\right]\ .$ (3)
@@ 4.4 TD learning using learned rewards
Despite not meeting the reward threshold, sub-optimal state-action tuples possess the potential to enhance the learned policy, augmenting its robustness against distribution shifts during deployment. This is due to the broader coverage of the state space within the auxiliary dataset, surpassing the limited span of the small expert dataset. In the absence of access to expert behavior in these states for direct learning, we employ a shortest path strategy to leverage these samples. More precisely, our objective on these states is to steer the policy toward states observed by the expert and subsequently imitate the expert’s behavior accordingly. To accomplish this, we use a TD3-style (Fujimoto & Gu, 2021) algorithm to learn a Q-function using the approximated rewards and then direct the policy to maximize the long-term return on these expert-unobserved states. The Q-function is learned as follows,
@@
$\min_{\pi}\underset{(s,a)\sim{\mathcal{D}}_{E}}{\mathbb{E}}\left[-\log\pi(a|s)\right]+\alpha\underset{(s,a)\sim{\mathcal{D}}_{O}}{\mathbb{E}}\left[-\log\pi(a|s)\cdot\tilde{r}(s,a)\cdot\mathbbm{1}[\tilde{r}>\tau]\right]+\beta\underset{s\sim\mathcal{D}_{E}\cup\mathcal{D}_{O}}{\mathbb{E}}\left[-Q(s,\pi(s))\right]\ .$ (6)
Here, $\beta$ is a hyperparameter which controls the contribution of the Q-loss. By maximizing this Q-function alongside the weighted BC objective, we incentivize the policy to: 1) act optimally in states where we have expert actions, and 2) guide the agent efficiently from expert-unobserved states to expert-observed states and act optimally subsequently.
The pseudo-code for the overall framework is presented in Algorithm 1.
2:  Train discriminator $d$ with non-negative PU learning following Eqn. 1
3:  for t=1 to T do
6:     Compute Q-function output values $Q(s,a)$ using sampled $(s,a)$ and $\mathcal{B}^{\pi}Q(s,a)$ using Eqn. 5
7:     Update $Q$ by minimizing the learning objective given in Eqn. 4
8:     if t mod $t_{freq}$ then
9:        Update $\pi$ by minimizing the learning objective in Eqn. 6
10:     end if
11:  end for
@@ 5 Experiments
  1. 1.
    How robust is ROIDA when the quality of the auxiliary data varies? (Sec. 5.3.1)
  2. 2.
    How does ROIDA compare to other methods as the size of the expert dataset changes? (Sec. 5.3.2)
@@ 5.1 Experimental setup
We conduct experiments on locomotion and manipulation tasks from the D4RL benchmark (Fu et al., 2020).
Locomotion
We use 4 MuJoCo environments for the locomotion tasks: hopper, halfcheetah, walker2d, and ant. For expert demonstrations in each environment, we use the corresponding dataset: hopper-expert-v2, halfcheetah-expert-v2, walker2d-expert-v2, and ant-expert-v2. For the sub-optimal demonstrations, we source trajectories from the respective random-v2 datasets. To create $\mathcal{D}_{E}$, we randomly sample 3, 5, or 7 trajectories per environment depending on the chosen setting. For $\mathcal{D}_{O}$, we create 3 settings per environment: 1000 randomly sampled sub-optimal trajectories plus another 0, 3, or 5 expert trajectories. This allows us to test our method’s ability to identify and leverage expert demonstrations within different mixes of sub-optimal and expert data.
Manipulation
We evaluate our method on 4 ADROIT manipulation tasks from using a simulated 24 DoF hand: pen twirling, hammering a nail, opening a door, and relocating a ball. For expert demonstrations, we sample 50 trajectories from pen-expert-v1, hammer-expert-v1, door-expert-v1, and relocate-expert-v1 respectively. For the sub-optimal demonstrations, we use 1000 trajectories from the datasets pen-cloned-v1, hammer-cloned-v1, door-cloned-v1, and relocate-cloned-v1, plus 0, 30, or 50 expert trajectories from the corresponding expert-v1 datasets.
Note that for all the evaluations, we report the average of the mean normalized score for the last 10 evaluations of training over 5 random seeds. Additional implementation details for our method can be found in Appendix A.2. We also present evaluations using the rliable (Agarwal et al., 2021) framework in Appendix A.3.
5.2 Baselines
We compare ROIDA against the following algorithms:
  • BC-exp: BC-exp denotes behavioral cloning solely on the expert dataset $\mathcal{D}_{E}$. Since $\mathcal{D}_{E}$ contains only a small number of expert demonstrations, training only on this data can lead to degraded performance at test time due to compounding errors caused by distribution shift.
  • BC-all: In this case, both $\mathcal{D}_{E}$ and $\mathcal{D}_{O}$ are used to learn the policy. Despite the large number of demonstrations, a significant portion of them are random or low quality. As a result, the learned policy tends to be sub-optimal due to the inclusion of poor demonstrations.
  • DWBC: DWBC (Xu et al., 2022) treats the auxiliary data as a mixture of expert and sub-optimal data, and utilizes PU learning to train a discriminator for weighted behavioral cloning. They perform a dual learning strategy where they alternately train the discriminator and the policy by taking the output of each model as an input to the other. Due to relying solely on BC, DWBC performs poorly when the number of expert transitions is low.
  • DemoDICE: DemoDICE (Kim et al., 2022) conducts state-action distribution matching over $\mathcal{D}_{E}$ and introduces a regularization constraint to ensure the learned policy remains close to the behavior policy of $\mathcal{D}_{O}$. It shares the same drawbacks as DWBC due to the second term, resulting in a suboptimal policy especially when $\mathcal{D}_{O}$ contains a large proportion of noisy data.
  • ORIL: ORIL (Zolna et al., 2020) first learns a reward function and then performs Critic-Regularized Regression (Wang et al., 2020) to learn the policy by enriching the data using different augmentation strategies.
@@ 5.2 Baselines
5.3 Results
5.3.1 Varying the quality of auxiliary data
Although DWBC achieves the second highest average performance, it degrades substantially compared to ROIDA when auxiliary data quality is poor (5/0 and 5/3). This stems from DWBC’s inability to extract useful information from poor quality unlabeled demos. In contrast, ROIDA can utilize the auxiliary data through TD learning, despite poor quality. Both DemoDICE and ORIL perform worse in all settings.
@@ 5.3.1 Varying the quality of auxiliary data
5.3.2 Varying the size of the expert dataset
In Table 2, we examine how varying the size of the expert dataset affects policy performance. For locomotion tasks, we fix the auxiliary dataset and set the number of expert trajectories in $\mathcal{D}_{O}$ to 5, while varying the number in $\mathcal{D}_{E}$ between 3, 5, and 7. For the Adroit tasks, we hold the number of expert trajectories in $\mathcal{D}_{O}$ at 50 and vary $\mathcal{D}_{E}$ from 50 to 70. ROIDA achieves comparable performance to DWBC, slightly outperforming in 7 out of 8 environments. As the amount of expert data in $\mathcal{D}_{E}$ increases, all the methods improve, as expected. In summary, ROIDA maintains competitive performance across varying expert dataset sizes.
5.4 Ablation study
In this section, we conduct experiments to analyze each component in ROIDA, and also benchmark the performance as the scale of the auxiliary dataset is varied.
First, we systematically remove each part of ROIDA and evaluate the resulting performance impact. The ablation results in Table 3 present the performance across three auxiliary dataset settings: 5/0, 5/3, 5/5 (denoting different auxiliary data qualities).
We begin by examining the impact of ablating the reward weighted BC component by setting $\alpha=0$. Performance drops indicate the importance of filtering the data using a learned reward model. The precise reward assignment enables ROIDA to discern the varying quality of transitions and utilize them accordingly.
Table 3: Ablation study on each module’s contribution to final policy performance on locomotion tasks.
Environment Setting Method
ROIDA w/o reward-weighted BC w/o modified reward w/o PU learning
Hopper 5 /  0 $84.63$ p m 16.01 $41.68$ p m 39.67 $78.07$ p m 11.00 $84.55$ p m 9.01
5 /  3 $86.66$ p m 21.94 $78.06$ p m 16.17 $92.06$ p m 7.10 $82.78$ p m 11.49
5 /  5 $88.45$ p m 8.46 $81.51$ p m 15.70 $84.73$ p m 11.26 $82.23$ p m 11.23
Avg. 86.58 p m 16.42 67.08 p m 26.34 84.96 p m 9.97 83.19 p m 10.64
HalfCheetah 5 /  0 $15.89$ p m 9.60 $9.42$ p m 1.48 $13.93$ p m 3.31 $15.74$ p m 8.63
5 /  3 $18.73$ p m 3.67 $14.93$ p m 8.61 $16.35$ p m 4.44 $14.53$ p m 3.02
5 /  5 $24.70$ p m 4.95 $18.22$ p m 4.61 $22.53$ p m 9.91 $16.68$ p m 7.26
Avg. 19.78 p m 6.58 14.19 p m 5.70 17.61 p m 6.55 15.65 p m 6.74
Walker2D 5 /  0 $108.73$ p m 0.28 $87.72$ p m 28.91 $104.65$ p m 7.79 $108.08$ p m 0.48
5 /  3 $108.52$ p m 0.21 $108.11$ p m 0.66 $104.77$ p m 6.83 $104.78$ p m 3.57
5 /  5 $108.794$ p m 0.09 $107.97$ p m 0.55 $105.56$ p m 4.81 $105.05$ p m 4.79
Avg. 108.68 p m 0.21 101.27 p m 16.70 104.99 p m 6.60 105.97 p m 3.46
Ant 5 /  0 $65.73$ p m 25.19 $57.42$ p m 20.89 $49.37$ p m 11.64 $70.05$ p m 18.71
5 /  3 $77.52$ p m 4.98 $64.40$ p m 11.52 $62.94$ p m 10.27 $69.77$ p m 10.31
5 /  5 $76.68$ p m 9.44 $58.95$ p m 17.68 $76.14$ p m 18.52 $69.08$ p m 11.06
Avg. 73.31 p m 15.79 60.25 p m 17.14 62.81 p m 13.95 69.63 p m 13.89
Next, we study the impact of the reward function formulation by directly using the discriminator output as the reward. The results demonstrate that our reward formulation based on the DICE methods improves policy performance compared to directly using the discriminator rewards.
We also examine the choice of PU learning (Eqn. 1) over binary classification for reward learning. PU learning accounts for the possibility of optimal transitions within the auxiliary dataset, whereas binary classification risks misclassifying all auxiliary samples as suboptimal, which could degrade performance.
When the auxiliary dataset contains only suboptimal transitions (5/0), both methods perform similarly. However, as the diversity of the auxiliary dataset increases (5/3 and 5/5), the performance of binary classification decreases. This drop occurs because binary classification incorrectly labels some optimal transitions as suboptimal, negatively impacting overall performance.
Setting Reward threshold
5 /  0 $83.20$ p m 24.58 $84.63$ p m 16.01 $70.24$ p m 27.75
5 /  3 $86.82$ p m 6.85 $86.66$ p m 21.94 $81.91$ p m 18.33
5 /  5 $88.23$ p m 11.42 $88.45$ p m 8.46 $90.70$ p m 6.88
Second, we study the impact of the reward threshold $\tau$. Our specific reward formulation leads to rewards in the range $[-2.19,2.19]$. Based on this, we decide to set the threshold $\tau=1$ to strike a balance between utilizing the “good” transitions in the auxiliary data and neglecting the poor quality transitions. To evaluate the impact of the threshold on the results, we vary on the Hopper environment as shown in Table 4.
Third, we change the scale of the auxiliary dataset and analyze the impact on the results. We achieve this by including 25%, 50% or 100% of the entire random set from D4RL (instead of the fixed 1000 trajectories in the previous experiments). As shown in Fig. 3, even with an increased number of suboptimal demonstrations, performance improves as more expert trajectories are added. This clearly highlights ROIDA’s ability to distinguish expert data even when the ratio of expert to noisy data is highly skewed. We also observe that adding more auxiliary data while keeping the number of expert trajectories fixed slightly improves performance, indicating ROIDA’s ability to extract information from suboptimal data.
6 Conclusion
We propose ROIDA, a simple yet effective framework for offline imitation that can maximize utilization of an unlabeled auxiliary dataset of unknown quality alongside a small set of expert demonstrations. Unlike previous methods that make assumptions about auxiliary dataset quality, ROIDA can seamlessly leverage uncurated, unlabeled offline datasets without relying on any quality assumptions. We demonstrate ROIDA’s efficacy on multiple manipulation and locomotion tasks, encompassing a wide variety of auxiliary dataset quality settings. The consistent performance gains over baselines validate ROIDA’s ability to unlock the full potential of heterogeneous offline datasets without relying on quality assumptions.
Limitations
Table 5: Performance on the Hopper task with ground-truth rewards.
Setting Method
ROIDA ROIDA w/ GT rewards
5 /  0 $84.63$ p m 16.01 $94.63$ p m 20.76
5 /  3 $86.66$ p m 21.94 $98.53$ p m 11.70
5 /  5 $88.45$ p m 8.46 $104.46$ p m 5.42
Avg. 86.58 p m 16.42 99.21 p m 14.11
In our particular framework, the reward estimation can be improved by an accurate choice of the hyperparameter $\eta$ by performing mixture proportion estimation (Zhu et al., 2023; Ramaswamy et al., 2016). However, this is beyond the scope of our work. In order to avoid any assumption about the auxiliary data in our work, we have chosen $\eta=0.5$ which is an unbiased estimate. We also provide additional results with $\eta=0.3$ and $\eta=0.7$ in Table 6. Here, we obtain better results when $\eta$ is closer ($\eta=0.3$ is closer than $\eta=0.7$) to the true ratio between expert and suboptimal demonstration in the auxiliary dataset (0.01 for setting 5/0, 0.12 for setting 5/3 and 0.19 for setting 5/5). Since this true ratio is unknown, estimating it would be a problem in its own right, which could then be combined with our method.
Table 6: Performance on the Hopper task with varying $\eta$.
Setting Method
5 /  0 $88.12$ p m 14.93 $84.63$ p m 16.01 $82.40$ p m 20.76
5 /  3 $90.42$ p m 18.84 $86.66$ p m 21.94 $84.85$ p m 7.13
5 /  5 $91.02$ p m 7.97 $88.45$ p m 8.46 $86.45$ p m 18.61
Avg. 89.85 p m 14.62 86.58 p m 16.42 84.57 p m 14.39
Additionally, the current implementation of ROIDA is designed for a single-task setting. An interesting avenue for future work is to extend our framework to multi-task or goal-conditioned settings, where the model is trained on datasets from different tasks or goals. In such a scenario, ROIDA could be used to learn a new, unseen, but related task based on these multi-task datasets. This extension would open up possibilities for more generalizable and flexible reinforcement learning applications across various tasks.
Broader impact statement
Acknowledgment
This work was partially supported by The National Science Foundation Award No. 2326309, National Institute for Food and Agriculture Award No. 2021-67022-33453 and the UC Multi-campus Research Programs Initiative.
Appendix A Appendix
A.1 Additional comparisons
We present additional comparisons in Table 7 with another recent unpublished algorithm, BCDP (Shao et al., 2023), which tailors an offline IL algorithm specifically for scenarios where the auxiliary offline dataset contains no expert data. While this design can enhance performance when the auxiliary dataset lacks expert demonstrations, it can lead to suboptimal performance as the proportion of expert data in the auxiliary dataset increases. The results indicate the ROIDA significantly outperforms BCDP on all the seven environments.
Method Environment
Hopper Walker2D Ant Pen Door Hammer Relocate
A.2 Experimental details
A.2.1 Implementation